OpenStudy (anonymous):

Adding rational expressions. I keep getting these wrong. x/x^2+6x+9 + 7/x^2+8x+15

Nnesha (nnesha):

\[\frac{ x }{ x^2+6x+7 } +\frac{ 7 }{ x^2+8x+15 }\] like this right ?

OpenStudy (anonymous):

that was dauntingly fast !

OpenStudy (anonymous):

x^2+6x+9 on the first one

OpenStudy (anonymous):

5(y+3)(y−1)/(y−3)2(y+1) my answer but I can't risk getting it wrong again.

OpenStudy (anonymous):

I've been doing something wrong with factoring on these. it's the correct answer but i'm assuming they want the full answer, not factored. but I have no clue

Nnesha (nnesha):

haha alright try to factor quadratic equations apply the AC (headphone method) find two number when you multiply them you should get product of AC and when you add or subtract them you should get middle term Ax^2+Bx+C=0 a=leading coefficient b=middle term c=constant ter

Nnesha (nnesha):

okay gimme a sec :)

Nnesha (nnesha):

so so what are the factors of x^2+6x+9 ??

Nnesha (nnesha):

so so what are the factors of x^2+8x+15??

OpenStudy (anonymous):

I have several of these to answer. will you help me with a few of them? I'm suppose to get 8 out of 10 correct and I was suppose to have all of this finished 22 minutes ago and I don't know if I submit this after i'm finish if it will even count. so i'm kinda panicked. haha I've done this test more than 10 times in the past week and still haven't figured out what i was doing wrong. decided to pay for owlbucks and have someone help me considering the deadline has already passed.

Nnesha (nnesha):

hmm okay good :)

Nnesha (nnesha):

do you know how to find factors quadratic equations ?

Nnesha (nnesha):

* of

OpenStudy (anonymous):

No, i don't

Nnesha (nnesha):

alright here is the easy method AC method (aka headphone method ) where you have to find two numbers when you multiply them you should get product of AC and when you add or subtract them you should get middle term  a= leading coefficient which is one b=middle term = 8 c=constant term = 15 so find two number when you multiply them you should get 15 but when you add or subtract them you should get 8